# Learning from Reward-Free Offline Data: A Case for Planning with Latent Dynamics Models

### v3

@@ 4.1 Two-Rooms Environment
We begin with a navigation task called Two-Rooms, featuring a point-mass agent. Each observation $x_{t}\in{\mathbb{R}}^{2\times 64\times 64}$ is a top-down view: the first channel encodes the agent, the second encodes the walls (Figure˜3). Actions $a\in{\mathbb{R}}^{2}$ denote the displacement vector of the agent position from one time step to the next, with a norm limit of $2.45$. The goal is to reach a randomly sampled state within 200 steps. See Section˜B.2 for more details. This environment allows for controlled data generation — ideal for efficient and thorough experimentation, while still not being so trivial that any method can solve it with even a little bit of data.
Offline data. To generate offline data, we place the agent in a random location within the environment, and execute a sequence of actions for $T$ steps, where $T$ denotes the episode length. The actions are generated by first picking a random direction, then using Von Mises distribution with concentration 5 to sample action directions. The step size is uniform from 0 to $2.45$. By default, the episodes’ length is $T=91$, and the dataset size is 3 million transitions.
@@ 4.2 What Methods Excel In-Distribution with a Large High-Quality Dataset?
To get the topline performance of the methods under optimal dataset conditions, we test them in a setting with abundant data, good state coverage, and trajectories long enough to traverse the two rooms. With 3 million transitions, corresponding to around 30,000 trajectories, all methods reach good performance in the goal-reaching task in Two-Rooms (Section˜3.3), with HIQL, HILP, and PLDM nearing 100% success rate.
@@ 4.3 What Method is the Most Sample-Efficient?
Figure 4: Testing the selected methods’ performance under different dataset constraints. Values and shaded regions are means and standard error over 3 seeds, respectively. Left: To test the importance of the dataset quality, we mix the random policy trajectories with good quality trajectories (see Figure˜3). As the amount of good quality data goes to 0, methods begin to fail, with PLDM and HILP being the most robust ones. Center: We measure methods’ performance when trained with different sequence lengths. We find that many goal-conditioned methods fail when train trajectories are short, which causes far-away goals to become out-of-distribution for the resulting policy. Right: We measure methods’ performance with datasets of varying sizes. We see that PLDM is the most sample efficient, and manages to get almost 80% success rate even with a few thousand transitions. See Appendix˜H for the analysis of statistical significance.
We investigate how different methods perform when the dataset size varies. While our ultimate goal is to have a method that can make use of large amounts of suboptimal offline data, this experiment helps distinguish which methods glean the most information from available data. We tried dataset sizes down to few thousand transitions. In Figure˜4 we see that the model-based method PLDM outperforms model-free methods when the data is scarce. In particular, HILP is more data-hungry than other model-free methods but achieves perfect performance with enough data.
Takeaway: PLDM is more sample-efficient than the model-free methods.
@@ 4.4 What Methods Can Stitch Suboptimal Trajectories?
Can we learn from short trajectories? We vary the episode length T during data generation to test whether methods can stitch together short training trajectories to reach long-horizon goals. In real-world scenarios, collecting long episodes is often difficult—–especially in open-ended environments—–so the ability to learn generalizable policies from short trajectories is crucial. The hardest scenario of the Two-Rooms environment requires the agent to navigate from the bottom left corner to the bottom right corner, and involves $\sim$ 90 steps, meaning that with short episode lengths such as 16, the goal is never observed. To succeed, the methods must stitch together multiple offline fragments.
We create datasets with episode length of 91, 64, 32, 16, keeping total transitions number at 3 million. Results in Figure˜4 (center) show that goal-conditioned methods fail when trained with shorter episodes. We hypothesize that because goal-conditioned methods sample most state and goal pairs from the same trajectory to train their policies, far-away goals become out of distribution for the resulting policy. Although randomly sampling goals from other trajectories during training does not improve generalization in our experiments, we hypothesize that data augmentation akin to the one outlined in [Ghu+24] can help. In contrast, HILP performs well by learning to follow directions in latent space, even from short episodes. Similarly, a model-based method such as PLDM can learn an accurate model from short trajectories and stitch together a plan during test time.
Can we learn from data with imperfect coverage? We artificially constrain trajectories to always stay within one room within the episode, and never pass through the door. Without the constraint, around 35% trajectories pass through the door. During evaluation, the agent still needs to go through the door to reach the goal state. This also reflects possible constraints in real-life scenarios, as the ability to stitch offline trajectories together is essential to efficiently learn from offline data. The results are shown in Section˜3.3. We see that HILP achieves perfect performance, while PLDM performance drops but still outperforms offline GCRL methods. As in the short-trajectory setting, GCRL methods fail. We hypothesize that HILP’s latent space structure enables effective stitching, while PLDM retains some performance due to the learned dynamics. Model-free GCRL fails because goals in different rooms are out-of-distribution for policies trained on same-room trajectories.
Takeaway: When solving a task requires ‘stitching’, HILP works well. The performance of PLDM drops slightly but is better than that of offline model-free GCRL.
@@ 4.5 What Methods Can Learn From Trajectories of a Random Policy?
We see that HILP and (model-based) PLDM outperform the goal-conditioned RL methods in this setting (Figure˜4). We hypothesize that because trajectories on average do not go far, the sampled state and goal pairs during training are close to each other, making faraway goals out of distribution and thus prevent GCRL from working effectively. In contrast, PLDM uses the data only to learn the dynamics model, and random policy trajectories are still suitable for the purpose. HILP learns the latent space and how to traverse it in various dimensions—even from random data.
Takeaway: When data quality is low, HILP and PLDM perform better than offline GCRL.
4.6 Which Methods Can Generalize to New Tasks?
To build effective general-purpose systems that learn from offline data, we need an algorithm that can generalize to different tasks. So far, we have evaluated on goal-reaching tasks. In this experiment, we test generalization to a different task: avoiding a chasing agent in the same environment. We compare PLDM and HILP, using models trained on optimal data from Section˜4.2 without any further training. In this task, the chaser follows an expert policy along the shortest path to the agent, and we vary its speed to adjust difficulty. The goal of the controlled agent is to avoid being caught. Goal-conditioned methods are excluded from evaluation, as they cannot avoid specific states by design. At each step, the agent observes the chaser’s state and selects actions to maintain distance. In PLDM, we invert the sign of the planning objective to maximize distance in the latent space to the goal state. In HILP, we invert the skill direction. We evaluate the success rate—defined as maintaining a distance $\geq$ 1.4 pixels over 100 steps. The results are shown in Figure˜5(b). We also plot average distance between the agents over time in Figure˜5(b). We see that PLDM performs better than HILP, and is able to evade the chaser more effectively by maintaining greater separation by episode end.
Takeaway: Planning with a latent dynamics model (PLDM) can generalize well to new tasks.
4.7 Extending to a Higher-Dimensional Control Environment
@@
We collect a dataset using a directional expert policy from [Par+24], generating 5M transitions by resampling a new direction every 10 steps and adding Gaussian noise with standard deviation 1.0 to each action. For evaluation, the quadruped is initialized in the bottom left or right corner, with the goal at the opposite diagonal. Each method is evaluated on 10 trials.
As in 4.4, we test the methods’ abilities to stitch short training trajectories by using datasets with trajectory lengths 25, 50, 100, 250, and 500 (during evaluation, the start and goal are $\approx 200$ steps apart). We show results in Fig.˜6. Consistent with the prior results, PLDM and HILP outperform GCRL baselines in trajectory stitching, achieving $100\%$ success rates while other methods fail with shorter trajectories.
Takeaway: Planning with a latent dynamics model (PLDM) maintains an advantage in solving a standard quadruped maze task, suggesting good generalization under higher control complexity.
4.8 What Methods Can Generalize to Unseen Environments?
In this experiment, we test whether methods can generalize to new environments—a key requirement for general-purpose agents, since collecting data for every scenario is infeasible. We introduce a new navigation environment with more complex dynamics and configurable layouts (Figure˜1). Building on top of Mujoco PointMaze, [TET12] layouts are generated by randomly permuting wall locations. Data is collected by randomly sampling actions at each step. Observation includes an RGB top-down view of the maze and agent velocity; actions are 2D accelerations. The goal is to reach a randomly sampled location (see Section˜B.2 for details).
To test generalization, we vary the number of training maze layouts (5, 10, 20, 40) and evaluate on held-out unseen layouts. For models trained on 5 maps, we further analyze how test-time performance degrades as layouts diverge from the training distribution. Figure˜7 shows that PLDM generalizes best – even when trained on just 5 maps – while other methods fail. As test layouts become more out-of-distribution, all methods except PLDM degrade in performance. We also evaluate all methods on a single fixed layout and observe 100% success rate across the board (Table˜4). Figure˜7 shows PLDM’s inferred plans and trajectories from different agents. We also investigate HILP’s failure to generalize in Appendix˜G, and show that HILP’s learned representation space successfully captures distance between states in mazes seen during training, but fails on unseen mazes.
Takeaway: PLDM enables better generalization to unseen environment variations than model-free methods.
@@ 5 Conclusion
In this work, we conducted a comprehensive study of existing methods for learning from offline data without rewards, spanning both reinforcement learning and optimal control, aiming to identify the most promising approaches for leveraging suboptimal trajectories. We focus on a set of navigation tasks that present unique challenges due to the need for spatial reasoning, generalization, and trajectory stitching. Our findings highlight HILP and PLDM as the strongest candidates, with PLDM demonstrating the best generalization to new environment layouts and tasks. Table˜1 shows a summary of our experimental results. Overall, we draw three main conclusions:
@@
1. Learning a latent dynamics model and using it to plan (PLDM) is robust to data quality, exhibits superior data efficiency, and enables generalization to new layouts and tasks; 2. Learning a well-structured latent-space (e.g., using HILP) enables trajectory stitching and robustness to data quality but is more data-hungry than other methods; 3. Model-free GCRL methods are a great choice when data is plentiful and high-quality.
Limitations. All experiments in this study were conducted in navigation environments, excluding robot manipulation or partially observable settings. While the fact that even in relatively simple settings we find that many methods fail in unexpected ways suggests that the conceptual insights into the effects of data quality on the investigated methods may hold more broadly, a more exhaustive study is needed.
Moving forward.
@@
Our findings indicate that learning and planning with latent dynamics models is a promising direction for building general autonomous agents. There are many interesting areas for exploration: 1) extending PLDM to more complex domains such as robotic manipulation and partially observable environments; 2) investigating improved dynamics learning methods to mitigate issues like accumulating prediction errors for tasks involving long-horizon reasoning [LPC22]; and 3) enhancing test-time efficiency, either by backpropagating gradients through the forward model [BXS20] or via amortized planning.
Appendix A Visualization of Plans and Trajectories for Diverse PointMaze
@@
Figure 7: Top: The training layouts used in the 5 layout setting. Middle: Trajectories of different agents navigating an unseen maze layout towards goal at test time. As the layouts become increasingly out-of-distribution, only PLDM consistently succeeds. Layouts can be represented as a 4x4 array, with each value being either a wall or empty space. The distribution shift is quantified by the minimum edit distance between a given test layout and the closest training layout. The top row corresponds to an in-distribution layout with a minimum edit distance of 0, and with each subsequent row, the minimum edit distance increases by 1 incrementally.
Appendix B Environments and Datasets
B.1 Two-Rooms Environment
@@
We build our own top-down navigation environment. It is implemented in PyTorch [Pas+19], and supports GPU acceleration. The environment does not model momentum, i.e. the agent does not have velocity, and is moved by the specified action vector at each step. When the action takes the agent through a wall, the agent is moved to the intersection point between the action vector and the wall. We generate the specified datasets and save them to disk for our experiments. The datasets generation takes under 30 minutes.
B.2 Diverse PointMaze
@@
Here, we build upon the Mujoco PointMaze environment [TET12], which contains a point mass agent with a 4D state vector $(\mathrm{global}\ x,\mathrm{global}\ y,v_{x},v_{y})$, where $v$ is the agent velocity. To allow our models to perceive the different maze layouts, we use as model input a top down view of the maze rendered as $(64,64,3)$ RGB image tensor instead of relying on $(\mathrm{global}\ x,\mathrm{global}\ y)$ directly.

### v4

@@ 4.1 Two-Rooms Environment
We begin with a navigation task called Two-Rooms, featuring a point-mass agent. Each observation $x_{t}\in{\mathbb{R}}^{2\times 64\times 64}$ is a top-down view: the first channel encodes the agent, the second the walls (Figure˜3). Actions $a\in{\mathbb{R}}^{2}$ denote the displacement vector of the agent position from one time step to the next, with a norm limit of $2.45$. The goal is to reach a randomly sampled state within 200 steps. See Section˜C.2 for more details. This environment allows for controlled data generation – ideal for efficient and thorough experimentation, while still not being too trivial.
Offline data. To generate offline data, we place the agent in a random location within the environment, and execute a sequence of actions for $T$ steps, where $T$ denotes the episode length. The actions are generated by first picking a random direction, then using Von Mises distribution with concentration 5 to sample action directions. The step size is uniform from 0 to $2.45$. Unless otherwise specified, the episodes’ length is $T=91$, and the total number of transitions in the data is 3 million.
@@ 4.2 What Methods Excel In-Distribution with a Large High-Quality Dataset?
To get the topline performance of the methods under optimal dataset conditions, we test them in a setting with abundant data, good state coverage, and good quality trajectories long enough to traverse the two rooms. With 3 million transitions, corresponding to around 30,000 trajectories, all methods reach good performance in the goal-reaching task in Two-Rooms (Section˜3.3), with HIQL, GCIQL, HILP, and PLDM nearing 100% success rate.
@@ 4.3 What Method is the Most Sample-Efficient?
Figure 4: Testing the selected methods’ performance under different dataset constraints. Values and shaded regions are means and standard error over 3 seeds, respectively. Left: To test the importance of the dataset quality, we mix the random policy trajectories with good quality trajectories (see Figure˜3). As the amount of good quality data goes to 0, methods begin to fail, with PLDM, GCIQL, and HILP being the most robust ones. Center: We measure methods’ performance when trained with different sequence lengths. We find that many goal-conditioned methods fail when train trajectories are short, which causes far-away goals to become out-of-distribution for the resulting policy. Right: We measure methods’ performance with datasets of varying sizes. We see that PLDM and GCIQL are the most sample efficient, and manage to get almost 80% success rate even with a few thousand transitions. See Appendix˜H for the analysis of statistical significance.
We investigate how different methods perform when the dataset size varies. While our ultimate goal is to have a method that can make use of a large amount of suboptimal offline data, this experiment serves to distinguish which methods can glean the most information from available data. We tried ranges of dataset sizes all the way down to a few thousand transitions. In Figure˜4, we see that the model-based PLDM and the model-free GCIQL outperform other model-free methods when data is scarce. In particular, HILP is more data-hungry than other model-free methods but achieves perfect performance with enough data.
Takeaway: PLDM and GCIQL are more sample-efficient than other methods.
@@ 4.4 What Methods Can Stitch Suboptimal Trajectories?
Can we learn from short trajectories? We vary the episode length T during data generation to test whether methods can stitch together short training trajectories to reach long-horizon goals. In real-world scenarios, collecting long episodes is often difficult—especially in open-ended environments—so the ability to learn generalizable policies from short trajectories is crucial. The hardest scenario of Two-Rooms requires the agent to navigate from the bottom left corner to the bottom right corner, and involves $\sim$ 90 steps, meaning that with short episode lengths such as 16, the goal is never observed. To succeed, methods must stitch together multiple offline trajectories.
We create datasets with episode length of 91, 64, 32, 16, keeping total transitions number at 3 million. Results in Figure˜4 (center) show that with the exception of GCIQL, goal-conditioned model-free methods struggle when trained with shorter episodes. We hypothesize that these methods are limited by dynamic programming on the short transitions, which can be sample-inefficient for stitching many short trajectories to reach far-away goals. In contrast, HILP performs well by learning to follow directions in latent space – even from short episodes. Similarly, PLDM can learn an accurate model from short trajectories and stitch together a plan during test time.
Can we learn from data with imperfect coverage? We artificially constrain trajectories to always stay within one room within the episode, and never pass through the door. Without the constraint, around 35% trajectories pass through the door. During evaluation, the agent still needs to go through the door to reach the goal state. This also reflects possible constraints in real-life scenarios, as the ability to stitch offline trajectories together is essential to efficiently learn from offline data. The results are shown in Section˜3.3. We see that HILP and GCIQL achieve perfect performance, while PLDM performance drops but is still higher than the rest of offline GCRL methods. We hypothesize that HILP’s latent space structure enables effective stitching, while PLDM retains some performance due to the learned dynamics. With the exception of GCIQL, other model-free GCRL methods fail to learn to compose trajectories across rooms.
Takeaway: When solving the task requires ‘stitching’, HILP and GCIQL work great. The performance of PLDM drops, but is better than that of most offline model-free GCRL methods.
@@ 4.5 What Methods Can Learn From Trajectories of a Random Policy?
We see that HILP, GCIQL and the model-based PLDM outperform the rest of goal-conditioned RL methods in this setting (Figure˜4). We hypothesize that because random trajectories on average do not go far, the sampled state and goal pairs during training are close to each other. As a result, faraway goals become out of distribution, and GCRL methods struggle with long-horizon tasks when TD learning fails to bridge distant trajectories. In contrast, PLDM uses the data only to learn the dynamics model, and random policy trajectories are still suitable for the purpose. HILP learns the latent space and how to traverse it in various dimensions – even from random data.
Takeaway: When the dataset quality is low, HILP, GCIQL and PLDM perform better than other offline GCRL methods.
4.6 What Methods Can Generalize to a New Task?
To build effective general purpose systems that learn from offline data, we need an algorithm that can generalize to different tasks. So far, we evaluated on goal-reaching tasks. In this experiment, we test generalization to a different task in an environment with the same dynamics: avoiding a chasing agent in the same environment. We compare PLDM and HILP, using models trained on optimal data from Section˜4.2 without any further training. In this task, the chaser follows an expert policy along the shortest path to the agent, and we vary its speed to adjust difficulty. The goal of the controlled agent is to avoid being caught. Goal-conditioned methods are excluded from evaluation, as they cannot avoid specific states by design. At each step, the agent observes the chaser’s state and selects actions to maintain distance. In PLDM, we invert the sign of the planning objective to maximize distance in the latent space to the goal state. In HILP, we invert the skill direction. We evaluate the success rate - defined as maintaining a distance $\geq$ 1.4 pixels over 100 steps. The results are shown in Figure˜5(b). We also plot average distance between the agents over time in Figure˜5(b). We see that PLDM performs better than HILP, and is able to evade the chaser more effectively by maintaining greater separation by episode end.
Takeaway: Assuming the environment dynamics remain fixed, PLDM can generalize to tasks other than goal-reaching simply by changing the planning objective.
4.7 Extending to a Higher-Dimensional Control Environment.
@@ 4.6 What Methods Can Generalize to a New Task?
We collect a dataset using a pretrained directional expert policy from Park et al. [51], generating 5M transitions by resampling a new direction every 10 steps and adding Gaussian noise with standard deviation 1.0 to each action. For evaluation, the quadruped is initialized at the bottom left or right corner, with the goal at the opposite diagonal. Each method is evaluated on 10 trials.
As in 4.4, we test the methods’ abilities to stitch short training trajectories by using datasets with trajectory lengths 25, 50, 100, 250, and 500 (during evaluation, the start and goal are approximately $200$ steps apart). PLDM, HIQL, and HILP outperform other GCRL baselines in trajectory stitching, achieving $100\%$ success rates while other methods fail with shorter trajectories. In contrast to the Two-Rooms experiments, HIQL outperforms GCIQL in this setting. We hypothesize that HIQL’s hierarchical structure may particularly benefit high-dimensional ant control: the low-level policy can manage fine-grained joint movements, while the high-level policy can govern overall navigation.
Takeaway: Planning with a latent dynamics model maintains good performance in a standard quadruped maze task, indicating promising generalization to higher control complexity.
4.8 What Methods Can Generalize to Unseen Environment Layouts?
In this experiment, we test whether methods can generalize to new obstacle configurations/layouts – a key requirement for general purpose agents, since collecting data for every scenario is infeasible. We introduce a new navigation environment with slightly more complex dynamics and configurable layouts (Figure˜2). Building on top of Mujoco PointMaze, [66] layouts are generated by randomly permuting wall locations. Data is collected by randomly sampling actions at each step. Observation includes an RGB top-down view of the maze and agent velocity; actions are 2D accelerations. The goal is to reach a randomly sampled location. See Section˜C.2 for details.
Figure 7: Left: Plans generated by PLDM at test time. Right: Actual agent trajectories for the tested methods. PLDM is the only method that reliably succeeds on held-out mazes.
To test generalization to new obstacle configurations, we vary the number of training maze layouts (5, 10, 20, 40) and evaluate on held-out unseen layouts. For models trained on 5 maps, we further analyze how test-time performance degrades as layouts diverge from the training distribution. Figure˜8 shows that PLDM generalizes best – even when trained on just 5 maps – while other methods fail. As test layouts become more out-of-distribution, all methods except PLDM degrade in performance. We also evaluate all methods on a single fixed layout and observe 100% success rate across the board (Table˜4). Figure˜7 and 8 show PLDM’s inferred plans and trajectories from different agents. We also investigate HILP’s failure to generalize in Appendix˜I, and show that HILP’s learned representation space successfully captures distance between states in mazes seen during training, but fails on unseen mazes.
Takeaway: The model-based approach enables better generalization to unseen obstacle layouts than model-free methods.
@@ 5 Conclusion
In this work, we conducted a comprehensive study of existing methods for learning from offline data without rewards, spanning both reinforcement learning and optimal control, aiming to identify the most promising approaches for leveraging suboptimal trajectories. We focus on a set of navigation tasks that present unique challenges due to the need for spatial reasoning, generalization to new layouts, and trajectory stitching. Our findings highlight HILP and PLDM as the strongest candidates, with PLDM demonstrating the best generalization to new obstacle layouts and to a state-avoidance task. We aggregate our experimental results in Table˜1. Overall, we draw three main conclusions:
@@
1. PLDM exhibits robustness to data quality, a high level of data efficiency, best-of-class generalization to new layouts, and excels at adapting to tasks beyond goal-reaching; 2. Learning a well-structured latent-space (e.g. using HILP) enables trajectory stitching and robustness to data quality, although it is more data-hungry than other methods; 3. Model-free GCRL methods are a great choice when the data is plentiful and good quality.
Future work.
@@
Our findings indicate that learning and planning with latent dynamics models is a promising direction for building general autonomous agents. There are many promising areas for exploration: 1) extending PLDM to more complex domains such as robotic manipulation and partially observable environments; 2) investigating improved dynamics learning methods to mitigate issues like accumulating prediction errors for tasks involving long-horizon reasoning [37]; and 3) improving test-time efficiency, either by backpropagating gradients through the forward model [8] or via amortized planning.
Acknowledgments
This work was supported through the NYU IT High Performance Computing resources, services, and staff expertise, by the Institute of Information & Communications Technology Planning & Evaluation (IITP) with a grant funded by the Ministry of Science and ICT (MSIT) of the Republic of Korea in connection with the Global AI Frontier Lab International Collaborative Research, by the Samsung Advanced Institute of Technology (under the project Next Generation Deep Learning: From Pattern Recognition to AI) and the National Science Foundation (under NSF Award 1922658), and in part by AFOSR grant FA9550-23-1-0139 "World Models and Autonomous Machine Intelligence" and by ONR MURI N00014-22-1-2773 "Self-Learning Perception through Interaction with the Real World".
Appendix
Appendix A Reproducibility
Code is available at https://github.com/vladisai/PLDM
An implementation of our method is provided in the GitHub repository above, which includes environments, data generation, training, and evaluation scripts. Detailed hyperparameters for our method and the baselines are provided in Appendix˜J.
Appendix B Visualization of Plans and Trajectories for Diverse PointMaze
@@
Figure 8: Top: The training layouts used in the 5 layout setting. Middle: Trajectories of different agents navigating an unseen maze layout towards goal at test time. As the layouts become increasingly out-of-distribution, only PLDN consistently succeeds. Layouts can be represented as a 4x4 array, with each value being either a wall or empty space. The distribution shift is quantified by the minimum edit distance between a given test layout and the closest training layout. The top row corresponds to an in-distribution layout with a minimum edit distance of 0, and with each subsequent row, the minimum edit distance increases by 1 incrementally.
Appendix C Environments and Datasets
C.1 Two-Rooms Environment
@@
We build our own top-down navigation environment. It is implemented in PyTorch [54], and supports GPU acceleration. The environment does not model momentum, i.e. the agent does not have velocity, and is moved by the specified action vector at each step. When the action takes the agent through a wall, the agent is moved to the intersection point between the action vector and the wall. We generate the specified datasets and save them to disk for our experiments. The datasets generation takes under 30 minutes.
C.2 Diverse PointMaze
@@
Here, we build upon the Mujoco PointMaze environment [66], which contains a point mass agent with a 4D state vector $(\mathrm{global}\ x,\mathrm{global}\ y,v_{x},v_{y})$, where $v$ is the agent velocity. To allow our models to perceive the different maze layouts, we use as model input a top down view of the maze rendered as $(64,64,3)$ RGB image tensor instead of relying on $(\mathrm{global}\ x,\mathrm{global}\ y)$ directly.